OpenStudy (anonymous):

If the perimeter of a square is 40 miles, what is its area?

OpenStudy (johnweldon1993):

Square = 4 equal sides.. If P = 40 40/4 = 10....each side is 10 miles long For area...do length * width.....but each side length is the same...10....so area = 10 * 10 = ...?